lithium borohydride crystallises in orthorhombic system with 4 molecules per unit cell.the unit cell dimensions are a=6.8A,b=4.4A,c=7.2A.if molar mass is 21.76,calculate density of cry

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shefali

File:Orthorhombic.png

volume of the unit cell = abc =6.8*4.4*7.2 =215.424 A= 2.154 *10-22 cc

each unit cell has 4 molecules  so mass of each unit cee  = 4*21.76/(6.023 *1023)

                                                                                =1.445 * 10-22  gm

so density =1.445 * 10-22  gm/2.154 *10-22 cc 

                    =.671 gm/cc

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin